OpenStudy (anonymous):

Can someone explain what function sequence is and how it works because my lesson does not explain well, and I need to know that information. Thanks in advance!

OpenStudy (studygurl14):

Hia Kaylee

OpenStudy (anonymous):

HEYA

OpenStudy (anonymous):

Can you help me

OpenStudy (perl):

maybe you could upload the worksheet

OpenStudy (anonymous):

remember perl the video

OpenStudy (studygurl14):

hi @perl I fanned you. even though I couldn't fan you by hovering over username, i went to your page and fanned you there

OpenStudy (perl):

yeah i skimmed the video

OpenStudy (anonymous):

lol

OpenStudy (perl):

thx @StudyGurl14

OpenStudy (perl):

this is like a matrix video about functions

OpenStudy (anonymous):

studygurl14, you know my name but I don't know yours

OpenStudy (studygurl14):

so....Kaylee, is this required for school as something you have to write up, or are you just wondering?

OpenStudy (perl):

@KayleeNeedsHelp do you get the idea of what a function is. its like a soda machine

OpenStudy (studygurl14):

function sequence is different from a regular function @perl

OpenStudy (perl):

you put in x (your money), the soda machine turns the money into a soda using some rule , and you get a soda

OpenStudy (perl):

study, the only difference is the inputs are 1,2,3..

OpenStudy (perl):

f(n)

OpenStudy (studygurl14):

ok true. don't mean to start another debate lol

OpenStudy (anonymous):

well in my lesson I have this about sequence in functions and it does not explain well

OpenStudy (perl):

can you ask a specific problem from your homework?

OpenStudy (anonymous):

perl, i know where you got that from......

OpenStudy (studygurl14):

yeah, show us a specific problem from your book and we'll explain it

OpenStudy (perl):

otherwise its kind of vague

OpenStudy (studygurl14):

yep ^

OpenStudy (perl):

Kaylee, i made up the soda example. but yeah, you can probably find the machine analogy online

OpenStudy (anonymous):

I don't have a book, I am homeschooled, connections academy

OpenStudy (studygurl14):

I'm with CA too and I have a math book...?

OpenStudy (anonymous):

I did'nt get one...

OpenStudy (perl):

she's with save the rainforest CA

OpenStudy (studygurl14):

mayeb that's why you aren't udnerstanding anything lol

OpenStudy (anonymous):

this is my 1st year

OpenStudy (anonymous):

i don't know if i am supposed to have one

OpenStudy (perl):

kaylee, you can read a book out of curiosity

OpenStudy (studygurl14):

the lessons have links to the online versionof textbooks if you didn't receive one. what math course are you taking?

OpenStudy (perl):

read a math book or math course

OpenStudy (anonymous):

algebra 1

OpenStudy (perl):

you can google function sequence

OpenStudy (anonymous):

something.....anything......

OpenStudy (studygurl14):

I'm taking Pre-calc...i never took alg 1at CA....like @perl said. just google it. the world wide web has some pretty good math websites

OpenStudy (anonymous):

well.... thanks anyway

OpenStudy (studygurl14):

kaylee give me a medal, i'll give perl a medal, and perl will give you a medal!

OpenStudy (perl):

see if you can find examples of function sequences online as in your lesson

OpenStudy (anonymous):

there

OpenStudy (anonymous):

give a medal to perl

OpenStudy (studygurl14):

i love doing medal exchanges

OpenStudy (perl):

@StudyGurl14 haha, that was cool

OpenStudy (anonymous):

cool=)

OpenStudy (anonymous):

you got an example?

OpenStudy (anonymous):

no,not really,sorry

OpenStudy (anonymous):

a sequence of functions usually looks like \[f_0,f_1,f_2,f_3,...\]

OpenStudy (perl):

i dont think she's doing a sequence of functions. she's doing a function sequence, so f(1),f(2),f(3)...

OpenStudy (anonymous):

you could say for example \[f_0=1, f_1=1+x, f_2=1+x+x^2,f_3=1+x+x^2+x^3...\]

OpenStudy (anonymous):

ooh sorry nvm

OpenStudy (perl):

but thats cool, never saw sequence function

OpenStudy (anonymous):

these are moderators, i thought they stayed checking everything was good?

OpenStudy (studygurl14):

moderators also help people with math. not just moderate

OpenStudy (anonymous):

are these like teachers too

OpenStudy (studygurl14):

sorta. they are even smarter than me! and that's hard to do. just kidding lol

OpenStudy (perl):

@satellite73 can you check me